Pierre Dupont just received a cash gift from his grandfather. He plans to invest in a five-year bond issued by Venice Corp. that pays an annual coupon rate of 4.43 percent. If the current market rate is 5.98 percent, what is the maximum amount Pierre should be willing to pay for this bond?
Answer #1

Annual coupon=1000*4.43%=44.3

Hence current value=Annual coupon*Present value of annuity factor(5.98%,5)+$1000*Present value of discounting factor(5.98%,5)

=44.3*4.214656645+$1000*0.747963532

=$934.67(Approx).

NOTE:

1.Present value of annuity=Annuity[1-(1+interest rate)^-time period]/rate

=44.3[1-(1.0598)^-5]/0.0598

=44.3*4.214656645

2.Present value of discounting factor=1000/1.0598^5

=1000*0.747963532